# A Theory of the NEPv Approach for Optimization On the Stiefel Manifold

### v2

@@ Appendix B Preliminary Lemmas
In this section, we collect a few results, some known and some likely new, that we need in our proofs. We will point out an earliest reference to each known one. Some likely appear before but we are not aware of any reference to. For completeness, we will provide proofs for those we cannot find references.
@@
Let $\phi\,:\,\mathfrak{D}\subseteq\mathbb{R}^{N}\to\mathbb{R}$ be convex and differentiable and let $\boldsymbol{x},\,\widetilde{\boldsymbol{x}}\in$ where $\mathfrak{D}$ is a convex domain in $\mathbb{R}^{N}$. If
@@
That $\operatorname{tr}(B^{\mu}C^{\nu})\leq\|B^{\mu}C^{\nu}\|_{\operatorname{tr}}$ is a corollary of Weyl’s majorant theorem [8, p.42]. Let $Q\in\mathbb{O}^{k\times k}$ such that $Q^{\operatorname{T}}B^{\mu}C^{\nu}\succeq 0$, which yields $\|B^{\mu}C^{\nu}\|_{\operatorname{tr}}=\operatorname{tr}(Q^{\operatorname{T}}B^{\mu}C^{\nu})$. Note that $B\succeq 0$ and thus $(Q^{\operatorname{T}}B^{\mu})^{\operatorname{T}}(Q^{\operatorname{T}}B^{\mu})=B^{2\mu}$, implying the singular values of $Q^{\operatorname{T}}B^{\mu}$ are given by $\{[\sigma_{i}(B)]^{\mu}\}_{i=1}^{k}$. Since $C\succeq 0$, the singular values of $C^{\nu}$ are $\{[\sigma_{i}(C)]^{\nu}\}_{i=1}^{k}$. Hence by Lemma B.4 and then by Lemma B.2, we get
@@
where $Q\in\mathbb{O}^{k\times k}$ such that $Q^{\operatorname{T}}B\succeq 0$.
@@
As a corollary of Lemma B.9, for any $H\in\mathbb{R}^{k\times k}$, if $H\not\succeq 0$ (which means either $H$ is not symmetric or $H$ is symmetric but indefinite or even negative semidefinite), then $\operatorname{tr}(H)<\|H\|_{\operatorname{tr}}$. Now let $H=U\Sigma V^{\operatorname{T}}$ be the SVD of $H$ [24] where $\Sigma\in\mathbb{R}^{k\times k}$ and set $Q=UV^{\operatorname{T}}$. Then $Q^{\operatorname{T}}H=V\Sigma V^{\operatorname{T}}\succeq 0$ and $\operatorname{tr}(Q^{\operatorname{T}}H)=\|H\|_{\operatorname{tr}}>\operatorname{tr}(H)$.
@@
Let $H\in\mathbb{R}^{n\times n}$ be symmetric and $P_{*}\in\mathbb{O}^{n\times k}$ whose column space ${\cal R}(P_{*})$ is the invariant subspace of $H$ associated with its $k$ largest eigenvalues. Suppose that $\lambda_{k}(H)-\lambda_{k+1}(H)>0$. Given $P\in\mathbb{O}^{n\times k}$, let
@@
We know that $\eta\geq 0$ by Fan’s trace minimization principle [20]. The second inequality in (B.3) is [35, Theorem 1] and can also be derived from some of the estimates in [68, Chapter 3] and by a minor modification to the proof of [39, Theorem 2.2]. It remains to show the first inequality in (B.3). Expand $P$ to $[P,P_{\bot}]\in\mathbb{O}^{n\times n}$. We find
@@
for any $\xi\in\mathbb{R}$ because $P_{\bot}^{\operatorname{T}}P=0$. But in what follows, we will take $\xi=[\lambda_{1}(H)+\lambda_{n}(H)]/2$. Next we expand $P_{*}$ to $[P_{*},P_{*\bot}]\in\mathbb{O}^{n\times n}$. Since the column space of $P_{*}$ is the invariant subspace of $H$ associated with the $k$ largest eigenvalues of $H$, we have
@@
    $\operatorname{tr}(P^{\operatorname{T}}B)\leq\|B\|_{\operatorname{tr}}$ for any $P\in\mathbb{O}^{n\times k}$;
@@
    $\operatorname{tr}(P^{\operatorname{T}}B)=\|B\|_{\operatorname{tr}}$ where $P\in\mathbb{O}^{n\times k}$ if and only if $B=P\Lambda$ with $\Lambda\succeq 0$;
@@
    $\max_{P\in\mathbb{O}^{n\times k}}\operatorname{tr}(P^{\operatorname{T}}B)=\|B\|_{\operatorname{tr}}$
@@
Lemma B.11 says that $\operatorname{tr}(P^{\operatorname{T}}B)$ is bounded above by $\|B\|_{\operatorname{tr}}$ always and the upper bound $\|B\|_{\operatorname{tr}}$ is achieved by any orthogonal polar factor $P_{*}$ of $B$ and also any maximizer of $\operatorname{tr}(P^{\operatorname{T}}B)$ over $P\in\mathbb{O}^{n\times k}$ is an orthogonal polar factor of $B$. For numerical computation, an orthogonal polar factor of $B$ can be constructed from the thin SVD $B=U\Sigma V^{\operatorname{T}}$ as $P_{*}=UV^{\operatorname{T}}$ of $B$. Conceivably, the closer $\operatorname{tr}(P^{\operatorname{T}}B)$ is to the upper bound, the closer $P$ approaches to an orthogonal polar factor of $B$. The results of the next lemma quantify the last statement.
@@
Let $B\in\mathbb{R}^{n\times k}$ and suppose $\operatorname{rank}(B)=k$. Let $P_{*}$ be the unique orthogonal polar factor of $B$. Given $P\in\mathbb{O}^{n\times k}$, let
@@
Now we expand $P$ to $[P,P_{\bot}]\in\mathbb{O}^{n\times n}$. We find
@@
Next, we show (B.7). The proof technique is borrowed from [76, Lemma 4.1] and [62, section 3.1]. Suppose now that $P^{\operatorname{T}}B\succ 0$. Following the proof of [76, Lemma 4.1], we can conclude that there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=P_{*}Q^{\operatorname{T}}$ satisfies
@@
Consider now item (c) for which ${\cal R}(P)={\cal R}(P_{*})$. Then $P=P_{*}W$ for some $W\in\mathbb{O}^{k\times k}$. Recall that $B=P_{*}\Lambda$ is the polar decomposition of $B$, and hence $\Lambda\succ 0$ and its eigenvalues are the singular values of $B$. Let $\Lambda=U\Gamma U^{\operatorname{T}}$ be the eigendecomposition of $\Lambda$ where $U\in\mathbb{O}^{k\times k}$ and $\Gamma=\operatorname{diag}(\sigma_{1}(B),\ldots,\sigma_{k}(B))$. Write $U^{\operatorname{T}}WU=[w_{ij}]\in\mathbb{O}^{k\times k}$. We know $-1\leq w_{ii}\leq 1$ for $1\leq i\leq N$. We still have by (B.9)
@@
by Lemma B.11, we get $f(P^{(i+1)})\geq f(P^{(i)})+\omega\eta_{i}\geq f(P^{(i)})$ by the NPDo Ansatz, implying that the sequence $\{f(P^{(i)})\}_{i=0}^{\infty}$ is monotonically increasing. Since $f(P)$ is assumed differentiable in $P$ and $\mathbb{O}^{n\times k}$ is compact, the sequence $\{f(P^{(i)})\}_{i=0}^{\infty}$ is uniformly bounded and hence convergent. This proves item (a).
@@
by design, we have $f(P^{(i+1)})\geq f(P^{(i)})+\omega\eta_{i}\geq f(P^{(i)})$ by the NEPv Ansatz, implying that the sequence $\{f(P^{(i)})\}_{i=0}^{\infty}$ is monotonically increasing. Since $f(P)$ is assumed differentiable in $P$ and $\mathbb{O}^{n\times k}$ is compact, sequence $\{f(P^{(i)})\}_{i=0}^{\infty}$ is bounded and hence convergent. This proves item (a).
@@
$\operatorname{tr}(\Omega_{*})=\operatorname{tr}(P_{*}^{\operatorname{T}}H(P_{*})P_{*})=\max_{P\in\mathbb{O}^{n\times k}}\operatorname{tr}(P^{\operatorname{T}}H(P_{*})P)=:S_{k}(H(P_{*})),$ (D.3)
@@
Since $S_{k}(H(P))$, $\operatorname{tr}(P^{\operatorname{T}}H(P)P)$, and $f(P)$ are continuous in $P\in\mathbb{O}^{n\times k}$, it follows from (D.2) that there is an $i_{0}\in\mathbb{I}$ such that
@@
$Z=\lim_{\mathbb{I}_{3}\ni i\to\infty}P^{(i+1)}\in\mathbb{O}^{n\times k}.$
@@ Appendix E The $M$-inner Product
To proceed, we let the Cholesky decomposition of $M$ be $M=R^{\operatorname{T}}R$ and set $Z=RP$. It can be verified that $P^{\operatorname{T}}MP=I_{k}$ if and only if $Z^{\operatorname{T}}Z=I_{k}$. Hence $Z=RP$ and $P=R^{-1}Z$ establish an one-one mapping between $Z\in\mathbb{O}^{n\times k}$ and $\{P\in\mathbb{R}^{n\times k}\,:\,P^{\operatorname{T}}MP=I_{k}\}$. Optimization problem (E.1) is equivalently turned into
@@
^$\widehat{\mathscr{H}}(Z):=\frac{\partial\hat{f}(Z)}{\partial Z}=Z\Lambda\quad\mbox{with}\quad\Lambda^{\operatorname{T}}=\Lambda\in\mathbb{R}^{k\times k},\quad Z\in\mathbb{O}^{n\times k}.$ (E.3)
@@
$\widehat{H}(Z)Z=Z\Omega,\quad\Omega^{\operatorname{T}}=\Omega,\quad Z\in\mathbb{O}^{n\times k},$ (E.7)
@@
Let $H(P)\in\mathbb{R}^{n\times n}$ be a symmetric matrix-valued function on $\mathbb{O}^{n\times k}$, satisfying
$H(P)P-\frac{\partial f(P)}{\partial P}=MP\,\mathscr{M}(P)\quad\mbox{for $P\in\mathbb{O}^{n\times k}$},$ (E.10)
where $\mathscr{M}(P)\in\mathbb{R}^{k\times k}$ is some matrix-valued function. $P\in\mathbb{O}^{n\times k}$ is a solution to the KKT condition (E.5) if and only if it is a solution to NEPv (E.9) and $\mathscr{M}(P)$ is symmetric.
@@
Throughout this section, $0\leq\theta\leq 1$, $B\succeq 0$ and $\operatorname{rank}(B)>n-k$ which ensures $\operatorname{tr}(P^{\operatorname{T}}BP)\geq s_{k}(B)>0$ for any $P\in\mathbb{O}^{n\times k}$.
@@
For $P,\widehat{P}\in\mathbb{O}^{n\times k}$, suppose either $\theta\in\{0,1\}$ or $\operatorname{tr}(P^{\operatorname{T}}AP+P^{\operatorname{T}}D)\geq 0$, and let $\widetilde{P}=\widehat{P}Q$ where $Q\in\mathbb{O}^{k\times k}$ is an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}D$. If (F.1) holds, then
@@
$\geq f(P)+\frac{1}{2}\left(\frac{s_{k}(B)}{S_{k}(B)}\right)^{\theta}\,\eta+[S_{k}(B)]^{-\theta}\,\Big{[}\|\widehat{P}^{\operatorname{T}}D\|_{\operatorname{tr}}-\operatorname{tr}(\widehat{P}^{\operatorname{T}}DP^{\operatorname{T}}\widehat{P})\Big{]},$ (F.6)
where $s_{k}(B)$ and $S_{k}(B)$ are the sum of the $k$ smallest eigenvalues and that of the $k$ largest eigenvalues of $B$, respectively.

### v4

@@ Appendix B Preliminary Lemmas
In this section, we collect a few results, some known and some likely new, that we need in our proofs. We will point out an earliest reference, if known, to each one. Some likely appear before but we are not aware of any reference to. For completeness, we will provide proofs for those we cannot find references.
@@
Let $\phi\,:\,\mathfrak{D}\subseteq\mathbb{R}^{N}\to\mathbb{R}$ be convex and differentiable and let $\boldsymbol{x},\,\widetilde{\boldsymbol{x}}\in\mathfrak{D}$, where $\mathfrak{D}$ is a convex domain in $\mathbb{R}^{N}$. If
@@
That $\operatorname{tr}(B^{\mu}C^{\nu})\leq\|B^{\mu}C^{\nu}\|_{\operatorname{tr}}$ is a corollary of Weyl’s majorant theorem [8, p.42]. Let $Q\in{\rm St}(k,k)$ such that $Q^{\operatorname{T}}B^{\mu}C^{\nu}\succeq 0$, which yields $\|B^{\mu}C^{\nu}\|_{\operatorname{tr}}=\operatorname{tr}(Q^{\operatorname{T}}B^{\mu}C^{\nu})$. Note that $B\succeq 0$ and thus $(Q^{\operatorname{T}}B^{\mu})^{\operatorname{T}}(Q^{\operatorname{T}}B^{\mu})=B^{2\mu}$, implying the singular values of $Q^{\operatorname{T}}B^{\mu}$ are given by $\{[\sigma_{i}(B)]^{\mu}\}_{i=1}^{k}$. Since $C\succeq 0$, the singular values of $C^{\nu}$ are $\{[\sigma_{i}(C)]^{\nu}\}_{i=1}^{k}$. Hence by Lemma B.4 and then by Lemma B.2, we get
@@
where $Q\in{\rm St}(k,k)$ such that $Q^{\operatorname{T}}B\succeq 0$.
@@
As a corollary of Lemma B.9, for any $H\in\mathbb{R}^{k\times k}$, if $H\not\succeq 0$ (which means either $H$ is not symmetric or $H$ is symmetric but indefinite or even negative semidefinite), then $\operatorname{tr}(H)<\|H\|_{\operatorname{tr}}$. Now let $H=U\Sigma V^{\operatorname{T}}$ be the SVD of $H$ [24] where $\Sigma\in\mathbb{R}^{k\times k}$ and set $Q=UV^{\operatorname{T}}$, an orthogonal polar factor of $H$. Then $Q^{\operatorname{T}}H=V\Sigma V^{\operatorname{T}}\succeq 0$ and $\operatorname{tr}(Q^{\operatorname{T}}H)=\|H\|_{\operatorname{tr}}>\operatorname{tr}(H)$.
@@
Let $H\in\mathbb{R}^{n\times n}$ be symmetric and $P_{*}\in{\rm St}(k,n)$ whose column space ${\cal R}(P_{*})$ is the invariant subspace of $H$ associated with its $k$ largest eigenvalues. Suppose that $\lambda_{k}(H)-\lambda_{k+1}(H)>0$. Given $P\in{\rm St}(k,n)$, let
@@
We know that $\eta\geq 0$ by Fan’s trace minimization principle [20]. The second inequality in (B.3) is [35, Theorem 1] and can also be derived from some of the estimates in [68, Chapter 3] and by a minor modification to the proof of [39, Theorem 2.2]. It remains to show the first inequality in (B.3). Expand $P$ to $[P,P_{\bot}]\in{\rm St}(n,n)$. We find
@@
for any $\xi\in\mathbb{R}$ because $P_{\bot}^{\operatorname{T}}P=0$. But in what follows, we will take $\xi=[\lambda_{1}(H)+\lambda_{n}(H)]/2$. Next we expand $P_{*}$ to $[P_{*},P_{*\bot}]\in{\rm St}(n,n)$. Since the column space of $P_{*}$ is the invariant subspace of $H$ associated with the $k$ largest eigenvalues of $H$, we have
@@
    $\operatorname{tr}(P^{\operatorname{T}}B)\leq\|B\|_{\operatorname{tr}}$ for any $P\in{\rm St}(k,n)$;
@@
    $\operatorname{tr}(P^{\operatorname{T}}B)=\|B\|_{\operatorname{tr}}$ where $P\in{\rm St}(k,n)$ if and only if $B=P\Lambda$ with $\Lambda\succeq 0$;
@@
    $\max_{P\in{\rm St}(k,n)}\operatorname{tr}(P^{\operatorname{T}}B)=\|B\|_{\operatorname{tr}}$
@@
Lemma B.11 says that $\operatorname{tr}(P^{\operatorname{T}}B)$ is bounded above by $\|B\|_{\operatorname{tr}}$ always and the upper bound $\|B\|_{\operatorname{tr}}$ is achieved by any orthogonal polar factor $P_{*}$ of $B$ and also any maximizer of $\operatorname{tr}(P^{\operatorname{T}}B)$ over $P\in{\rm St}(k,n)$ is an orthogonal polar factor of $B$. For numerical computation, an orthogonal polar factor of $B$ can be constructed from the thin SVD $B=U\Sigma V^{\operatorname{T}}$ as $P_{*}=UV^{\operatorname{T}}$ of $B$. Conceivably, the closer $\operatorname{tr}(P^{\operatorname{T}}B)$ is to the upper bound, the closer $P$ approaches to an orthogonal polar factor of $B$. The results of the next lemma quantify the last statement.
@@
Let $B\in\mathbb{R}^{n\times k}$ and suppose $\operatorname{rank}(B)=k$. Let $P_{*}$ be the unique orthogonal polar factor of $B$. Given $P\in{\rm St}(k,n)$, let
@@
Now we expand $P$ to $[P,P_{\bot}]\in{\rm St}(n,n)$. We find
@@
Next, we show (B.7). The proof technique is borrowed from [76, Lemma 4.1] and [62, section 3.1]. Suppose now that $P^{\operatorname{T}}B\succ 0$. Following the proof of [76, Lemma 4.1], we can conclude that there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=P_{*}Q^{\operatorname{T}}$ satisfies
@@
Consider now item (c) for which ${\cal R}(P)={\cal R}(P_{*})$. Then $P=P_{*}W$ for some $W\in{\rm St}(k,k)$. Recall that $B=P_{*}\Lambda$ is the polar decomposition of $B$, and hence $\Lambda\succ 0$ and its eigenvalues are the singular values of $B$. Let $\Lambda=U\Gamma U^{\operatorname{T}}$ be the eigendecomposition of $\Lambda$ where $U\in{\rm St}(k,k)$ and $\Gamma=\operatorname{diag}(\sigma_{1}(B),\ldots,\sigma_{k}(B))$. Write $U^{\operatorname{T}}WU=[w_{ij}]\in{\rm St}(k,k)$. We know $-1\leq w_{ii}\leq 1$ for $1\leq i\leq N$. We still have by (B.9)
@@
by Lemma B.11, we get $f(P^{(i+1)})\geq f(P^{(i)})+\omega\eta_{i}\geq f(P^{(i)})$ by the NPDo Ansatz, implying that the sequence $\{f(P^{(i)})\}_{i=0}^{\infty}$ is monotonically increasing. Since $f(P)$ is assumed differentiable in $P$ and ${\rm St}(k,n)$ is compact, the sequence $\{f(P^{(i)})\}_{i=0}^{\infty}$ is uniformly bounded and hence convergent. This proves item (a).
@@
by design, we have $f(P^{(i+1)})\geq f(P^{(i)})+\omega\eta_{i}\geq f(P^{(i)})$ by the NEPv Ansatz, implying that the sequence $\{f(P^{(i)})\}_{i=0}^{\infty}$ is monotonically increasing. Since $f(P)$ is assumed differentiable in $P$ and ${\rm St}(k,n)$ is compact, sequence $\{f(P^{(i)})\}_{i=0}^{\infty}$ is bounded and hence convergent. This proves item (a).
@@
$\operatorname{tr}(\Omega_{*})=\operatorname{tr}(P_{*}^{\operatorname{T}}H(P_{*})P_{*})=\max_{P\in{\rm St}(k,n)}\operatorname{tr}(P^{\operatorname{T}}H(P_{*})P)=:S_{k}(H(P_{*})),$ (D.3)
@@
Since $S_{k}(H(P))$, $\operatorname{tr}(P^{\operatorname{T}}H(P)P)$, and $f(P)$ are continuous in $P\in{\rm St}(k,n)$, it follows from (D.2) that there is an $i_{0}\in\mathbb{I}$ such that
@@
$Z=\lim_{\mathbb{I}_{3}\ni i\to\infty}P^{(i+1)}\in{\rm St}(k,n).$
@@ Appendix E The $M$-inner Product
To proceed, we let the Cholesky decomposition of $M$ be $M=R^{\operatorname{T}}R$ and set $Z=RP$. It can be verified that $P^{\operatorname{T}}MP=I_{k}$ if and only if $Z^{\operatorname{T}}Z=I_{k}$. Hence $Z=RP$ and $P=R^{-1}Z$ establish an one-one mapping between $Z\in{\rm St}(k,n)$ and $\{P\in\mathbb{R}^{n\times k}\,:\,P^{\operatorname{T}}MP=I_{k}\}$. Optimization problem (E.1) is equivalently turned into
@@
^$\widehat{\mathscr{H}}(Z):=\frac{\partial\hat{f}(Z)}{\partial Z}=Z\Lambda\quad\mbox{with}\quad\Lambda^{\operatorname{T}}=\Lambda\in\mathbb{R}^{k\times k},\quad Z\in{\rm St}(k,n).$ (E.3)
@@
$\widehat{H}(Z)Z=Z\Omega,\quad\Omega^{\operatorname{T}}=\Omega,\quad Z\in{\rm St}(k,n),$ (E.7)
@@
Let $H(P)\in\mathbb{R}^{n\times n}$ be a symmetric matrix-valued function on ${\rm St}(k,n)$, satisfying
$H(P)P-\frac{\partial f(P)}{\partial P}=MP\,\mathscr{M}(P)\quad\mbox{for $P\in{\rm St}(k,n)$},$ (E.10)
where $\mathscr{M}(P)\in\mathbb{R}^{k\times k}$ is some matrix-valued function. $P\in{\rm St}(k,n)$ is a solution to the KKT condition (E.5) if and only if it is a solution to NEPv (E.9) and $\mathscr{M}(P)$ is symmetric.
@@
Throughout this section, $0\leq\theta\leq 1$, $B\succeq 0$ and $\operatorname{rank}(B)>n-k$ which ensures $\operatorname{tr}(P^{\operatorname{T}}BP)\geq s_{k}(B)>0$ for any $P\in{\rm St}(k,n)$.
@@
For $P,\widehat{P}\in{\rm St}(k,n)$, suppose either $\theta\in\{0,1\}$ or $\operatorname{tr}(P^{\operatorname{T}}AP+P^{\operatorname{T}}D)\geq 0$, and let $\widetilde{P}=\widehat{P}Q$ where $Q\in{\rm St}(k,k)$ is an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}D$. If (F.1) holds, then
@@
$\geq f(P)+\omega\,\eta+[S_{k}(B)]^{-\theta}\,\Big{[}\|\widehat{P}^{\operatorname{T}}D\|_{\operatorname{tr}}-\operatorname{tr}(\widehat{P}^{\operatorname{T}}DP^{\operatorname{T}}\widehat{P})\Big{]},$ (F.6)
where
$\omega=\begin{cases}\frac{1}{2}\left(\frac{s_{k}(B)}{S_{k}(B)}\right)^{\theta},\quad&\mbox{if $\eta\geq 0$},\\ \frac{1}{2}\left(\frac{S_{k}(B)}{s_{k}(B)}\right)^{\theta},\quad&\mbox{if $\eta<0$},\end{cases}$
$s_{k}(B)$ and $S_{k}(B)$ are the sum of the $k$ smallest eigenvalues and that of the $k$ largest eigenvalues of $B$, respectively.